The length of a side of the base of a square pyramid is 6m, the slant height is 27m. Find the surface area of the pyramid.
Appreciate any help in advance, I'm confused.

Answer by lwsshak3(11628)

The length of a side of the base of a square pyramid is 6m, the slant height is 27m. Find the surface area of the pyramid
..
The pyramid has 4 identical surfaces plus the square base.
1 surface area=1/2*base*height=1/2*6*27=81
square base=6*6=36
Total surface area=4*81+36=360 sq m